OpenStudy (anonymous):

Tom receives allowance every week. He spends three-fourths of his allowance on food. Last week, Tom spent nine dollars on food. What is Tom’s weekly allowance? Write the verbal model, define the variable, write the equation and find the solution. Show your work.

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (texaschic101):

3/4 of his allowance on food....he spends $9 on food...so $9 is 3/4 of his allowance. (3/4) / 9 = (4/4)/x -- (3/4 to $9 = 4/4 to x) cross multiply 3/4x = 9(4/4) 3/4x = 9 x = 9/(3/4) x = 9 * 4/3 x = 36/3 x = 12 check.. 3/4 of 12 = 9 3/4 * 12 = 9 36/4 = 9 9 = 9 (correct) Tom's weekly allowance is $12

OpenStudy (anonymous):

thank youu so much that deserves a medal

OpenStudy (texaschic101):

aww....thanks :)

OpenStudy (anonymous):

no thank you